a continuous random variable T has the following probability density function.

f(t) = { 32t2/128, -4≤t<4

0, otherwise


a. find E(t)

b. find Var(T)

c. Var(2T+3)


1
Expert's answer
2021-06-03T15:44:00-0400

"f(t)=\begin{cases}
 32t^2/128 &\text{if } -4\leq t<4 \\
 0 &\text{otherwise } 
\end{cases}"



First of all we have to check whether this function is probability density function

So for pdf "\int_{-\infty}^\infty f(t)dx=1"


"\int_{-\infty}^\infty f(t)dt=\int_{-4}^4 \dfrac{32t^2}{128}dt"

"= \dfrac{32}{128}\times\dfrac{1}{3}[t^3]_{-4}^4\\\ \\= \dfrac{32}{128}\times\dfrac{1}{3}[4^3-(-4)^3]\\\ \\=\dfrac{32}{128}\times \dfrac{1}{3}\times128\\\ \\=\dfrac{32}{3}\neq 1"


So, it is not a probability density function.


(a) "E(t)=\int_{-\infty}^\infty tf(t)dt=\int_{-4}^4 \dfrac{32t^3}{128}dt=\dfrac{32}{128}\times\dfrac{1}{4}[t^4]_{-4}^4=\dfrac{8}{128}[256-256]=0"


Hence, "E(t)=0"



(b) "Var(t)" "=E[(t-\mu)^2]= \int_{-\infty}^\infty (t-0)^2f(t)dt= \int_{-\infty}^\infty t^2f(t)dt=\int_{-4}^4 \dfrac{32}{128}t^4"


"\implies \dfrac{32}{128}\times \dfrac{1}{5}[t^5]_{-4}^4=\dfrac{32}{640}[1024-(-1024)]=\dfrac{32\times 2048}{640}=102.4"



(c) "Var(2t+3)=E[(2t+3-\mu)^2]= \int_{-\infty}^\infty (2t+3-0)^2f(t)dt= \int_{-\infty}^\infty (2t+3)^2f(t)dt"


"\implies \int_{-\infty}^\infty (4t^2+12t+9)f(t)dt=4 \int_{-\infty}^\infty t^2f(t)dt+12 \int_{-\infty}^\infty tf(t)dt+9 \int_{-\infty}^\infty f(t)dt"


We have ,

"\int_{-\infty}^\infty f(t)dt=32/3\\\ \\ \int_{-\infty}^\infty tf(t)dt=0\\\ \\ \int_{-\infty}^\infty t^2f(t)dt=102.4"


So,


"\implies \int_{-\infty}^\infty (4t^2+12t+9)f(t)dt=4 \int_{-\infty}^\infty t^2f(t)dt+12 \int_{-\infty}^\infty tf(t)dt+9 \int_{-\infty}^\infty f(t)dt\\\ \\\implies 4(102.4)+12\cdot(0)+9\times\dfrac{32}{3}=409.6+0+96=505.6"